PI: Gurvits
Title: VSOP test

Test: Jan 1997
Frequency: 1.3 cm
Obs time (UT): 13 EU 10:00(13/01)-14:40(13/01)
Stations: EfMhGoHnKpPtOv
Processor: Socc
EU Mk3 Tapes/St.: Europe 2 (thick), VLBA 1 (thin)


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Mon Jan 13 08:28:50 EST 1997) Minor failures
two scans after tape change messed up, only half on tape. Otherwise OK. Rapidly melting snow and sunshine means gain changes during experiment. No phase cal. Confirm 2dB slope in some 16MHz passbands, but see no problems with VC9 (implicated in vt6).

Metsahovi (Ari Mujunen/Kaj Wiik at Tue Jan 14 03:03:14 EST 1997) Success
Everything went smoothly, there was no need to edit the procedures this time (except at one point but that is considered to be fault in our setup and to be corrected ASAP). Weather was cloudy but no snow/rain. No formatter clock jumps, no stucking headstacks, no pointing computer problems...